OpenStudy (justin_lee):

In representing the eigenstates of angular momentum, eigenfunctions can be found using the ladder operators listed below: L+ = Lx + iLy L- = Lx - iLy In all the sources that I read, these ladder operators seem to be defined. Is there a reason to define the ladder operators this way? In other words, where do these two ladder operators come from? My guess is that since L^2 and Lz are hermitian, the difference L^2 - Lz^2 must be real. Since L^2 is defined as: L^2 = Lx^2 + Ly^2 + Lz^2 we can subtract Lz^2 from both sides to get L^2 - Lz^2 = Lx^2 + Ly^2 which has to equal a real value.

OpenStudy (justin_lee):

And L+ and L- are the attempted factorizations of Lx^2 and Ly^2 in the above equation.

OpenStudy (anonymous):

L+ and L- are not terms of factoring Lx^2 + Ly^2. Lx and Ly are operators, and do not commute to allow this to be so. \[(L_x + i L_y)(L_x - iL_y) = L_x^2 + L_y^2 +iL_yL_x - iL_xL_y\] but \[[L_x,L_y] = i \hbar L_z\] \[(L_x + i L_y)(L_x - iL_y) = L_x^2 + L_y^2 - i[L_x,L_y] = L_x^2+L_y^2+\hbar L_z\] As for why they are of this form, it's so they show the required behavior of ladder operators. If [O,L] = cL, where c is a scalar, allow |n> to be an eigenvector of O with eigenvalue n: \[O|n> = n|n>\]Then \[OL|n> = (OL+LO-LO)|n> = (LO + [O,L])|n> = (n+c)L|n>\] If c is real and positive, L is a raising operator. If c is real and negative, we call it a lowering operator. Often, it's convenient to also require that \[L_+^\dagger = L_-\]

OpenStudy (anonymous):

Small correction: \[L_+^\dagger = L_-\] Is a consequence of the operator being hermitian, along with flipping the sign on the scalar, c.

OpenStudy (michele_laino):

Such operators are introduced, since they are very useful. I can show the physical meaning of those operators: let's consider an eigenstate of L^2 and L_z: \[\begin{gathered} {L^2}\left| {a,b} \right\rangle = a\left| {a,b} \right\rangle \hfill \\ {L_z}\left| {a,b} \right\rangle = b\left| {a,b} \right\rangle \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

then we can write: \[\Large \begin{gathered} {L_z}{L_ + }\left| {a,b} \right\rangle = \left( {\left[ {{L_z},{L_ + }} \right] + {L_ + }{L_z}} \right)\left| {a,b} \right\rangle = \hfill \\ \hfill \\ = \left( {b + \hbar } \right)\left( {{L_z}\left| {a,b} \right\rangle } \right) \hfill \\ \end{gathered} \] since: \[\Large \left[ {{L_z},{L_ + }} \right] = \hbar {L_z}\]. Similarly for operator: \[\Large {L_ - }\]

OpenStudy (michele_laino):

oops.. I have made an error, here are the right formulas: \[\Large \begin{gathered} {L_z}{L_ + }\left| {a,b} \right\rangle = \left( {\left[ {{L_z},{L_ + }} \right] + {L_ + }{L_z}} \right)\left| {a,b} \right\rangle = \hfill \\ = \left( {b + h} \right){L_ + }\left| {a,b} \right\rangle \hfill \\ \end{gathered} \] where: \[\Large \begin{gathered} \left[ {{L_z},{L_ + }} \right] = {L_z}{L_ + } - {L_ + }{L_z}, \hfill \\ \hfill \\ \left[ {{L_z},{L_ + }} \right] = \hbar {L_ + } \hfill \\ \end{gathered} \] similarly goes, for the operator \[\Large {L_ - }\]